OpenStudy (loser66):

calculate \[\int {sin(2t)*e^{\frac{t}{2}}}dt\]

OpenStudy (loser66):

\[\int {sin(2t)*e^{\frac{t}{2}}}dt\]My prof uses something like \[cos (2t)+isin(2t)=e^{2it}\\\int{e^{2it}*e^{\frac{t}{2}}}dt\] \[=\frac{e^{(\frac{1}{2}+2i)t}}{\frac{1}{2}+2i}\] \[=\frac{e^{\frac{t}{2}}(cos (2t)+isin(2t)}{\frac{1}{2}+2i}\] Does anyone give me a link to figure out what the stuff is? Please.

OpenStudy (anonymous):

have you tried integration by parts?

OpenStudy (anonymous):

here this might explain a few things

OpenStudy (loser66):

one more thing likes \[Asin u \pm Bcos u=\sqrt {A^2+B^2}sin(u \pm arctan\frac{B}{A}\]

OpenStudy (anonymous):

from what i can tell your prof. uses the trig identities for this stuff

OpenStudy (anonymous):

I don't really see how Euler's formula helps here as you aren't using complex variables..I would just say use integration by parts.

OpenStudy (loser66):

Please, explain me

OpenStudy (anonymous):

For the last one, check out the derivation of the formula: http://www.mathsisfun.com/money/compound-interest-derivation.html I think that textbook you're using just assumes you've come across it at an earlier point and is now recalling it.

OpenStudy (loser66):

thank you, I will read it. The first part I type above, I attach, please check whether my question is stupid?

OpenStudy (anonymous):

I think you're right to question your prof's method here. This is not the way I would go about the integral. I'm still trying to decipher the work.

OpenStudy (loser66):

Thanks for response . just the beginning of the course. haia.....

OpenStudy (loser66):

Thanks for the link. It's new to me and helpful. Never took accounting before---> don't know the formulae